OpenStudy (valirium):

What is the solution of the following system? 3x + 3y = 10 -9x -9y = -30

jhonyy9 (jhonyy9):

rewrite the second one -3(3x+3y) = -3(10) and divide both sides by -3 so what will get ?

jhonyy9 (jhonyy9):

Syed your opinion about this ?

OpenStudy (valirium):

I'm confused

OpenStudy (htnsbs):

@jhonyy9 is the second one -9?

OpenStudy (mathmale):

There are several ways in which you could approach this problem. One would be to DIVIDE all 3 terms in the 2nd equation by 3. Write the first equation and then below it write the second equation. Combine like terms. What do you get?

OpenStudy (valirium):

@mathmale 3x + 3y = 10 -3x - 3 = -10 Am I right?

OpenStudy (valirium):

-3x - 3y = -10

jhonyy9 (jhonyy9):

yes correct

OpenStudy (mathmale):

Now you have 2 equations that are alike except for sign. Add them together. What result do you get? what does that result tell you?

OpenStudy (valirium):

@mathmale no solution?